# InfoLossQA: Characterizing and Recovering Information Loss in Text Simplification

### v1

@@ 3.3 Analysis of InfoLossQA
To characterize the information typically lost, we categorize questions according to 10 pragmatic function types defined by Cao and Wang (2021). To automate this analysis, we designed a few-shot prompting classifier based on the annotation guidelines of the authors. [^6] Prompting gpt-4-0613 with few-shot in-domain examples achieves an F1 score of 0.87 on a validation set. See Appendix A for details. Table 1 shows the question distribution across types, their definition and examples.
Most questions are of type procedural, requesting specific details about how the study was conducted (34.3%). This aligns with the earlier observation that most information loss occurs in Methods sections. The second most common type of questions seek more concrete definitions for overgeneralized concepts (25.7%), such as the health condition or clinical procedure. Comparing deletions and oversimplifications, we observe a similar distribution, with the exception of questions asking for the consequences (e.g., experimental results) or causes (e.g., what motivated the study) of an event being more prevalent among deletions.
In addition to the domain-agnostic question taxonomy above, we also analyze the QA pairs through the lens of the Paper Plain Key Question Index which was designed to convey the most important elements of an RCT August et al. (2023). We manually classify 120 QA pairs (40 per annotator) according to the 8 categories of the question index. Aligned with the above findings, the information which is most likely lost are about methodological detail (i.e., study protocol, analysis tools, population; 53% of QA pairs), followed by results (18% of QA pairs) and the goals of the trial (11% of QA pairs). The most critical information (usual and new treatment) are preserved. See Table A.4 for a detailed breakdown and examples.
@@
First, we consider the extent to which annotators highlight the same information loss spans. We calculate three measures on the evidence spans $E$ of the original text: (1) strict span-level F1, where boundaries and type (deletion vs. oversimplification) have to match exactly, (2) partial span-level F1 where two spans are considered a match if their token-level intersection over union exceeds a threshold DeYoung et al. (2020a), [^7] We tokenize on whitespace and use a threshold of $t=0.5$. and (3) sentence-level Krippendorff’s $\alpha$ where we project span-level annotations to a binary label indicating if a sentence has at least one span Goyal et al. (2022). As shown in Table 2, the sentence-level $\alpha$ shows that the annotators do tend to pick up information loss localized in the same evidence sentence to a moderate degree Artstein and Poesio (2008). However, the precise evidence span (in smaller linguistic units) is more subjective, as indicated by the lower span-level scores.
@@
Second, we analyze the similarity of questions $Q$ as a function of whether their underlying evidence spans E have complete, partial, or no overlap. Generally, questions become more similar when the underlying evidence is the same (Figure 3). However, subjectivity remains: Some questions are different even though the evidence is the same. Similar observations were made for human-generated QUDs Ko et al. (2022).
Granularity (Measure) Del. Over. All
Span, exact (F1) 0.20 0.13 0.17
Span, partial (F1) 0.34 0.34 0.34
Sentence (Krippendorff’s $\alpha$) 0.50
Table 2: Evidence span ($E$) overlap between annotators. On a span-level, overlap is reported separately for deletions (del) and oversimplifications (over).
Figure 3: Similarity of annotators’ questions ($Q$) by whether their evidence spans ($E$) have no/partial/complete overlap. We calculate cosine similarity over sentence embeddings Reimers and Gurevych (2019). Cosine greater 1 is an artifact of the kernel density estimate.
4 Methods for Automatically Identifying Information Loss
@@
No system in existing literature is directly applicable to solve all parts of the InfoLossQA task. We propose and evaluate two systems: (1) End-to-end LLM prompting, contrasting several strong open-source and commercial LLMs; (2) A pipeline that combines ideas from fact extraction, NLI, and question-generation into a full system.
4.1 LLM Prompting
@@
We design a one-shot prompt that contains a detailed specification of the task, closely resembling the manual annotation guidelines. [^8] Compared to a zero-shot prompt, we found that even one example substantially improves performance. For a fair comparison across models, we do not include more than one example as they would exhaust the context windows of some models under investigation. We ask all models to localize information loss instances, provide a short motivation for why this instance is considered information loss, akin to Chain-of-Thought prompting Wei et al. (2022), and finally to generate the QA pairs. We benchmark three LLMs: Mistral-7B, Llama2-70B, and GPT-4.999Specifically, we use gpt-4-0613, llama-2-70b-chat, and mistral-7b-instruct-v0.1. All prompts were manually tuned for each model, though this was a heuristic process and better prompts may exist. See Appendix A for prompts and decoding parameters.
4.2 NLI Pipeline
To contrast with end-to-end LLMs, we implement a three-step pipeline: (i) fact extraction, (ii) entailment prediction, (iii) QA generation for facts with a neutral relation.
NLI aims to determine whether a “hypothesis” can be inferred from a “premise,” classifying the hypothesis as entailed, contradictory or neutral. We use this framework to reason about the entailment of information in the original text (= hypothesis) with respect to its simplification (= premise). Intuitively, a neutral relation should hold when there is at least one piece of information in the original text which cannot be inferred from its simplification, i.e., an instance of information loss. In contrast, entailment indicates that all information is conveyed, and contradictions should not occur if the simplification is faithful.
A key challenge is the granularity of relation classification. The predominant approach in NLI is to classify at the sentence-level Bowman et al. (2015) with some extensions to document-level Schuster et al. (2022); Yin et al. (2021). However, for our purposes sentence-level is too coarse. Information loss instances can be as fine-grained as a single word or phrase, and one sentence may contain multiple distinct instances. To address this, we follow Stacey et al. (2023), who proposed entailment classification at the level of atomic facts. Reasoning over fine-grained facts is an established strategy, for example, in summarization evaluation Nenkova and Passonneau (2004); Liu et al. (2023), but a bottleneck is obtaining such facts. Stacey et al. (2023) showed that LLMs can effectively decompose sentences into shorter sentences, each encapsulating an atomic fact.
@@ 4.2 NLI Pipeline
We reconcile these ideas into a pipeline as follows. First, we use an LLM to decompose each sentence in the original text $X_{\text{src}}$ into a list of atomic facts. Then, a standard NLI model Laurer et al. (2024) classifies the entailment of each fact against the simple text $X_{\text{tgt}}$. Finally, we prompt an LLM to generate QA pairs given $X_{\text{src}}$, $X_{\text{tgt}}$, and the list of neutral/missing facts. The sentence that generated the fact serves as evidence $E$ [^10] We do not implement localization $R$ in the simple text, but note that interpretable NLI models could provide such a localization Stacey et al. (2022); Rodriguez et al. (2023). See Appendix A for prompts and decoding parameters.
@@ 5 Evaluation Framework
We next propose a comprehensive framework to evaluate automatic systems for InfoLossQA. This framework considers: (i) Correctness of the questions and answers (i.e., the QA pairs); (ii) Linguistic suitability for user-facing settings, and; (iii) Coverage of human QAs. All evaluation is done manually by trained annotators. We outline the evaluation criteria below and provide full annotator guidelines and examples in Appendix C.
5.1 Evaluation Criteria
Angle 1: Correctness of generated QA pairs.
@@ 5.1 Evaluation Criteria
The second criterion, Answer Accuracy, states that questions should be correctly answered by the generated/extracted answers ($A$s). We evaluate this in three parts. First, Accuracy - Evidence (3-point) captures how well the highlighted evidence $E$ answers $Q$. Second, Accuracy - Answer (3-point) evaluates how well $A$ answers $Q$. Finally, we ask if $A$ contains any Hallucinations (binary).
Angle 2: Linguistic suitability.
@@
As discussed in Section 2, we aim for QA pairs that are pragmatically felicitous, and written in a style that is comprehensible for readers. We propose evaluating this through four linguistic criteria. First, we evaluate Givenness (binary), closely following the constraints of QUD Riester et al. (2018); Wu et al. (2023): A question should only contain concepts that are either directly mentioned in the question context, inferable, or generally known. We define the question context as the entire simple text $X_{\text{tgt}}$ for QAs on Deletions, and everything up to and including the question localization $R$ for QAs on Oversimplifications. [^11] With this definition of question context we intend to enable interactions where readers can choose to inspect Oversimplification QAs as they read the text, for example, visualized as expandable highlights Fok et al. (2023).
Second, when a question seeks clarification about an oversimplified concept, there should be a rationale span $R$ in the simplified text. We evaluate this with the Rationale Localization criterion (3-point: good, unrelated, missing). Third, we evaluate if $A$ contains any Jargon (4-point). Finally, answers should be Standalone (binary): they should not contain any unresolved coreferences or acronyms that would require reading the original text. This is akin to the requirements for decontextualization Choi et al. (2021).
Angle 3: Recall of human QA.
The third angle evaluates if models identify similar information loss as humans, i.e., the recall of reference QAs in InfoLossQA. The challenge in this evaluation is that both humans and models produce a set of QAs without a strict 1-to-1 correspondence. It is possible that a QA on one side semantically entails zero, one, or more QAs on the other side. In a similar setting, Chen et al. (2022) directly compare the two sets, asking annotators to mark all QA pairs which are not entailed by the other set. However, we found this setup too cognitively demanding with typical sets of more than five QAs.
Therefore, we opt for a pairwise setting, comparing each reference QA with one model QA at a time, and marking the reference as either fully recalled, partially recalled or not recalled. As comparing all possible pairs is intractable, we follow a paraphrase mining approach to get candidate matches Wright et al. (2022). We use sentence embeddings Reimers and Gurevych (2019) to determine the similarity of two QA pairs. Afterwards, we establish a threshold by annotating 400 pairs equally sampled from 20 bins in the similarity range of $[0,1]$ and observe the first bin with a notable number of recalled QA (here: $T=0.65$). Pairs with similarity $\leq T$ are set to not recalled. [^12] For our dataset of 1,000 reference QAs across 104 documents and predictions by four models, this process reduced the total necessary comparisons by 93% from 33,825 to 2,446.
5.2 Annotation Protocol
We sample 100 QA pairs per model for angle 1+2, and 50% of the documents for angle 3.131313For angle 1+2, we take a stratified sample to preserve the relative frequencies at which models generate QAs per RCT section. As the generations by Mistral-7B and Llama2-70B do not always follow the specified output format, we only sample from QA pairs which could be completely parsed. See Table A.3 for a detailed analysis of generation errors. All samples were evaluated by two experts from the pool described in Section 3.2. The total evaluation effort was 64 hours for angle 1+2, and 20 hours for angle 3, with additional training hours. We observe a generally high agreement with slightly lower scores for Jargon and Answerability (Table 3).
Criterion Kappa
$Q$ is Answerable w/ $X_{\text{src}}$ 0.61
$Q$ is Answerable w/ $X_{\text{tgt}}$ 0.50
Accuracy - Evidence ($E$) 0.71
Accuracy - Answer ($A$) 0.68
Hallucinations ($A$) 0.89
Givenness ($Q$) 0.79
Rationale Localization ($R$) 0.80
Jargon ($A$) 0.57
Standalone ($A$) 0.69
Recall of human QA 0.70
Table 3: Inter-annotator agreement for manual evaluation: Randolph’s Kappa Randolph (2005).
@@ 5.2 Annotation Protocol
Model
Yes
${}^{\downarrow}$
${}^{\downarrow}$
No
Vague
${}^{\downarrow}$
Correct
${}^{\downarrow}$
${}^{\downarrow}$
Correct
${}^{\downarrow}$
${}^{\downarrow}$
Mistral-7B
72.5
19.6
7.8
32.4
30.4
37.3
74.5
12.7
12.7
84.8
12.3
2.9
Llama2-70B
83.0
9.7
7.3
21.4
32.5
46.1
77.7
11.7
10.7
85.9
10.7
3.4
GPT-4
87.5
12.5
0.0 58.5 33.0 8.5 98.0 2.0 0.0 94.0 6.0 0.0
NLI Pipeline 89.7 7.8
2.5
49.5
25.0
25.5
77.5
4.4
18.1
77.9
20.6
1.5
Table 4: Correctness of generated QA pairs (Angle 1) as percent of ratings given by two independent annotators over 100 QA pairs per model. ${}^{\downarrow}$ = lower is better; Multi. = Multiple valid answers; Part. = Partially complete answer.
Givenness ($Q$) Rationale Localization ($R$) Jargon ($A$) Standalone ($A$) Hallu. ($A$)
Model % Good Good ${}^{\downarrow}$ ${}^{\downarrow}$ (++) (+) (${}^{\downarrow}$ (${}^{\downarrow}$ Good % Good
Mistral-7B 74.5 52.0 5.9 42.2 21.6 9.3 13.2 55.9 84.3 96.1
Llama2-70B 75.2 34.5 5.3 60.2 26.2 7.8 17.5 48.5 79.6 92.2
GPT-4 82.0 52.0 1.5 46.5 15.0 24.0 11.0 50.0 90.5 99.0
NLI Pipeline 58.3 17.6 0.0 82.4 28.9 29.9 12.7 28.4 86.3 98.0
Table 5: Linguistic suitability of generated QA pairs (Angle 2) and hallucinations (hallu). We provide the percent of ratings given by two independent annotators over 100 QA pairs per model. ${}^{\downarrow}$ = lower is better.

### v2

@@ 3.3 Analysis of InfoLossQA
Question Type % All % O % D Examples
Procedural. Asking for the procedures, tools, or methods by which a certain outcome is achieved. 34.3 34.1 34.7 • How did they measure the patients’ opioid medication needs?
• How did the study control for bias?
Concept. Asking for a definition of an event or a concept. 25.7 28.1 18.9 • What kind of hip surgery were patients undergoing? • What type of mental illnesses are being studied?
Extent. Asking for the extent or quantity of an event or a concept. 17.0 19.2 10.8 • How many patients were in each group? • How much lurasidone was given to the patients in the trial?
Comparison. Asking for comparison among multiple events or concepts. 8.3 8.9 6.6 • How much did the abnormal blood vessels reduce in group 1 compared to group 2? • How did pre-meal insulin affect platelet activation compared to the placebo?
Consequence. Asking for the consequences or results of an event. 5.7 3.0 13.5 • What else does increased blood flow do? • What was the correlation between measures for total daily calcium intake?
Cause. Asking for the cause or reason for an event or a concept. 4.3 1.6 12.0 • What was a motivation behind the study? • Why are soy products being looked at in this study?
Example. Asking for example(s) or instance(s) of an event or a concept. 3.4 4.3 0.8 • What are the other brain functions that were tested? • What medication is typically used for outpatients?
Disjunction. Asking for the true one given multiple events or concepts, where comparison among options is not needed. 0.6 0.5 0.8 • Which one of the three components did this group’s supplement contain? • Which gel saw the increase in beneficial microbes after 24 hours?
Verification. Asking for the truthfulness of an event or a concept. 0.7 0.3 1.9 • Were there any other changes in blood sugar level observed in the study? • Were both eyes of each patient tested?
Judgmental. Asking for the opinions of the answerer’s own. Not observed.
Granularity (Measure) Del. Over. All
Span, exact (F1) 0.20 0.13 0.17
Span, partial (F1) 0.34 0.34 0.34
Sentence (Krippendorff’s $\alpha$) 0.50
Table 2: Evidence span ($E$) overlap between annotators. Del = Deletions; Over = Oversimplification.
@@
To characterize the information that is typically lost, we categorize questions according to the 10 pragmatic function types defined by Cao and Wang (2021) using a few-shot prompting classifier. [^7] Prompting gpt-4-0613 with annotation guidelines of Cao and Wang (2021) and few-shot examples achieves an F1 score of 0.87 on a validation set (Table A.1). Table 1 shows the question distribution across types, their definition and examples. Most questions are of type procedural, requesting specific details about how the study was conducted (34.3%). This aligns with the earlier observation that most information loss occurs in Methods sections. The second most common type of questions seek more concrete definitions for overgeneralized concepts (25.7%), such as the health condition or clinical procedure. Comparing deletions and oversimplifications, we observe a similar distribution, with the exception of questions asking for the consequences (e.g., experimental results) or causes (e.g., what motivated the study) of an event being more prevalent among deletions.
Section A.1 also provides a domain-specific classification of QAs in terms of frequently asked questions about RCTs August et al. (2023).
@@
First, we analyze to what extent annotators highlight the same evidence spans $E$. We calculate three measures: (1) strict span-level F1, where boundaries and type (deletion vs. oversimplification) have to match exactly, (2) partial span-level F1 where two spans are considered a match if their token-level intersection over union exceeds a threshold DeYoung et al. (2020a), [^8] We tokenize on whitespace and use a threshold of $t=0.5$. and (3) sentence-level Krippendorff’s $\alpha$ where we project span-level annotations to a binary label indicating if a sentence has at least one span Goyal et al. (2022). The sentence-level $\alpha$ shows that the annotators do tend to pick up information loss localized in the same evidence sentence to a moderate degree Artstein and Poesio (2008). However, the precise evidence span (in smaller linguistic units) is more subjective, as indicated by the lower span-level scores (Table 2).
@@
Second, we analyze the similarity of questions as a function of whether their underlying evidence spans $E$ have complete, partial, or no overlap. We use cosine similarity over question embeddings Reimers and Gurevych (2019). Generally, questions become more similar when the evidence overlaps (Figure 3(b)). However, subjectivity remains: Some questions are different although the evidence is the same. Similar observations were made for QUD generation Ko et al. (2022).
4 Automatic Information Loss Detection
@@
No system in literature is directly applicable to solve all parts of InfoLossQA. Therefore, we propose and evaluate two systems: (1) End-to-end prompting, contrasting several strong LLMs; (2) An NLI pipeline. We describe the methods below and provide detailed parameters in Table A.1.
Method 1: LLM prompting.
@@
We design a one-shot prompt that specifies the task, closely resembling the manual annotation guidelines. We ask all models to localize information loss instances, to provide a short motivation for why it is considered information loss, akin to Chain-of-Thought prompting Wei et al. (2022), and finally to generate the QA pairs. We benchmark three LLMs: Mistral-7B, Llama2-70B, and GPT-4.999Specifically, we use gpt-4-0613, llama-2-70b-chat, and mistral-7b-instruct-v0.1. All prompts were manually tuned for each model, though this was a heuristic process and better prompts may exist.
Method 2: NLI pipeline.
To contrast with end-to-end LLMs, we implement a three-step pipeline: (i) fact extraction, (ii) entailment classification, (iii) QA generation for facts with a neutral relation. NLI assesses if a “hypothesis” is inferable from a “premise,” categorizing it as entailed, contradictory, or neutral. We treat the original text $X_{\text{src}}$ as hypothesis, and the simplification $X_{\text{tgt}}$ as premise. Intuitively, a neutral relation indicates information loss, where at least one piece of information in $X_{\text{src}}$ cannot be inferred from $X_{\text{tgt}}$. In contrast, entailment indicates faithful conveyance of all information.
@@
Information loss can be as granular as a word or phrase, and one sentence may contain multiple distinct instances. Therefore, we adopt NLI over fine-grained facts Stacey et al. (2023). Formally, we decompose each sentence in $X_{\text{src}}$ into atomic facts using an LLM Stacey et al. (2023). Then, a standard NLI model Laurer et al. (2024) classifies entailment of each fact against $X_{\text{tgt}}$. Finally, we prompt an LLM to generate QA pairs given $X_{\text{src}}$, $X_{\text{tgt}}$, and the list of neutral facts. The sentence that generated the fact serves as evidence $E$ [^10] We do not implement localization $R$ in the simple text, but note that interpretable NLI models could provide such a localization Stacey et al. (2022); Rodriguez et al. (2023).
@@ 5 Evaluation Framework
We next propose a comprehensive framework to evaluate automatic systems for InfoLossQA. All evaluation is done manually by trained annotators. We describe the evaluation criteria below and Appendix C provides full annotator guidelines.
Aspect 1: Correctness of generated QA pairs.
@@
The second criterion, Answer Accuracy, states that questions should be correctly answered by the extracted and generated answers. We evaluate this in three parts. First, Accuracy - Evidence (3-point) captures how well the highlighted evidence $E$ answers $Q$. Second, Accuracy - Answer (3-point) evaluates how well $A$ answers $Q$. Finally, we ask if $A$ contains any Hallucinations (binary).
Aspect 2: Linguistic suitability.
@@
We aim for QAs that are pragmatically felicitous and comprehensible for readers (cf. Section 2). We propose evaluating this through four linguistic criteria. First, we evaluate Givenness (binary), closely following the constraints of QUD Riester et al. (2018); Wu et al. (2023): A question should only contain concepts that are directly mentioned in the question context, inferable, or generally known. We define the question context as the entire simple text $X_{\text{tgt}}$ for QAs on Deletions, and everything up to and including the rationale span $R$ for QAs on Oversimplifications. [^11] With this definition we intend to enable interactions where users can inspect oversimplification QAs as they read the text, e.g., visualized as expandable highlights Fok et al. (2023). Second, when a question seeks clarification about an oversimplified concept, there should be a rationale $R$ in $X_{\text{tgt}}$ as an opportunity to show users that the question addresses said vagueness (Rationale, 3-point). Third, we evaluate if $A$ contains any unexplained Jargon (4-point). Finally, Standalone (binary) states that answers must be decontextualized Choi et al. (2021), i.e., they should not contain unresolved coreferences or acronyms that would require reading the original text.
Aspect 3: Recall of human QA.
Lastly, we evaluate if models identify similar information loss as the references in InfoLossQA. A challenge in this evaluation is that both humans and models produce a set of QAs without a strict 1-to-1 correspondence. We found that directly comparing two question sets is too cognitively demanding with typical sets of more than five QAs. Therefore, we opt for a pairwise setting, comparing each reference QA with one model QA at a time, and marking the reference as either fully recalled, partially recalled or not recalled. As comparing all possible pairs is intractable, we exclude unlikely matches following a paraphrase mining approach Wright et al. (2022). We report the detailed procedure to estimate a minimum similarity threshold in Section A.3.
Evaluation process.
We sample 100 QAs per model for aspect 1+2, and 50% of the documents for aspect 3. All samples were independently evaluated by two of the experts described in Section 3.2. Evaluation took 64 hours for aspect 1+2, and 20 hours for aspect 3, with additional training hours. We measure Randolph’s Kappa Randolph (2005) for inter-evaluator agreement and observe a substantial average agreement of $\kappa=0.70$ (Table 5). Section A.3 provides details on sampling and exclusion of invalid LLM generations.
@@
Model Yes Multi. No No Vague Yes Correct Part. Wrong Correct Part. Wrong
Mistral-7B 72.5 19.6 7.8 32.4 30.4 37.3 74.5 12.7 12.7 84.8 12.3 2.9
Llama2-70B 83.0 9.7 7.3 21.4 32.5 46.1 77.7 11.7 10.7 85.9 10.7 3.4
GPT-4 87.5 12.5 0.0 58.5 33.0 8.5 98.0 2.0 0.0 94.0 6.0 0.0
NLI Pipeline 89.7 7.8 2.5 49.5 25.0 25.5 77.5 4.4 18.1 77.9 20.6 1.5
Table 3: Correctness of generated QA pairs as percent of ratings given by two independent annotators over 100 QA pairs per model. Multi. = Multiple valid answers; Part. = Partially complete answer; = lower is better.
Given. ($Q$) Rationale ($R$) Jargon ($A$) Stand. ($A$) Hallu. ($A$) Recall
Model % Good Good Unrel. Miss. (++) (+) (-) (–) % Good % Good Full Partial No
Mistral-7B 74.5 52.0 5.9 42.2 21.6 9.3 13.2 55.9 84.3 96.1 14.2 4.1 81.8
Llama2-70B 75.2 34.5 5.3 60.2 26.2 7.8 17.5 48.5 79.6 92.2 16.2 4.4 79.3
GPT-4 82.0 52.0 1.5 46.5 15.0 24.0 11.0 50.0 90.5 99.0 28.1 3.4 68.4
NLI Pipeline 58.3 17.6 0.0 82.4 28.9 29.9 12.7 28.4 86.3 98.0 22.9 20.0 57.2
Table 4: Linguistic suitability of generated QA pairs and recall of human QAs. Given. = Givenness; Stand. = Standalone; Hallu. = Hallucinations; Unrel. = Unrelated; Miss = Missing. = lower is better.